40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Sep 28 21:17:41 2012, Koji, Update, IOO, MC WFS adjustment MC_WFS_12h.png
    Reply  Mon Oct 1 11:08:25 2012, Jenne, Update, IOO, MC WFS adjustment 
Message ID: 7452     Entry time: Fri Sep 28 21:17:41 2012     Reply to this: 7455
Author: Koji 
Type: Update 
Category: IOO 
Subject: MC WFS adjustment 

MC WFS was fixed. Now it is running constantly with the autolocker.

Found a bug in the IOO screen: All of the 6 WFS signal indicators is liked to the same info (C1:IOO-MC1_PIT_OUTPUT).
Fix this, Jenne! Baaaaagghhhhh! 


What I did:

1. C1:IOO-MC_RFPD_DCMON indicator was saturating. "HOPR" of this entry was  set to 5 by running the following command:

ezcawrite C1:IOO-MC_RFPD_DCMON.HOPR 5

2. Scan MC2 spot position by using /opt/rtcds/caltech/c1/scripts/MC/moveMC2 scripts.
or the adjustment, C1:SUS-MC2_ASCPIT_EXC and C1:SUS-MC2_ASCYAW_EXC were excited with 300cnt at 12Hz and 10Hz, respectively.
The corresponding peaks (i.e. ANgle to length coupling) in C1:IOO-MC_F were monitored on DTT and adjusted so that the peaks are approximately nulled.

3. moveMC2 scripts are not perect to keep the maximum of the transmission. So, the alignment was adjusted with MC1 and MC3.

4. Repeated 2 and 3 until the alignment converges.

5. Once I got satisfied with the MC2 spot position, I went to the MC2 table and aligned the steering mirror before the QPD.

6. As these actions above moves the REFL beam, I went to the MC REFL path and adjusted the MC REFL PD position and the MC WFS spot positions.

7. Checked if the alignment is still good. The MC REFL is 0.50~0.51. Pretty good.

8. Run /opt/rtcds/caltech/c1/scripts/MC/WFS/WFS_FilterBank_offsets to register the current WFS offset etc.

9. At this point, MC WFS started working fine. I also confirmed the autolocker worked with this setting.

--------

Checked how the things are going in the morning. There were several unlocks. But the autolocker and WFS kept the cavity lcoked again.
Very good.

Some power fluctuation of ~1% is observed in the MC trans. I checked the PMC trans and found it is also fluctuating by 1% in a coherent way.
So I judge the WFS itself is fine. (See attached)

 

 

Attachment 1: MC_WFS_12h.png  35 kB  Uploaded Sat Sep 29 08:45:48 2012  | Hide | Hide all
MC_WFS_12h.png
ELOG V3.1.3-